### v2

# A Comprehensive Survey on Continual Learning in Generative Models

Abstract
The rapid advancement of generative models has enabled modern AI systems to comprehend and produce highly sophisticated content, even achieving human-level performance in specific domains. However, these models remain fundamentally constrained by catastrophic forgetting — a persistent challenge where adapting to new tasks typically leads to significant degradation in performance on previously learned tasks. To address this practical limitation, numerous approaches have been proposed to enhance the adaptability and scalability of generative models in real-world applications. In this work, we present a comprehensive survey of continual learning methods for mainstream generative models, encompassing large language models, multimodal large language models, vision-language action models, and diffusion models. Drawing inspiration from the memory mechanisms of the human brain, we systematically categorize these approaches into three paradigms: architecture-based, regularization-based, and replay-based methods, while elucidating their underlying methodologies and motivations. We further analyze continual learning setups for different generative models, including training objectives, benchmarks, and core backbones, offering deeper insights into the field. The project page of this paper is available at https://github.com/Ghy0501/Awesome-Continual-Learning-in-Generative-Models.
Index Terms:
Continual Learning, Generative Models, LLM, MLLM, VLA, Diffusion Models, Catastrophic Forgetting
1 Introduction
The evolution of artificial intelligence is undergoing a paradigm shift from “understanding the world” to “creating the world”. Traditional discriminative models, such as classification networks [1, 2, 3, 4, 5, 6] and object detectors [7, 8, 9, 10, 11], have achieved remarkable success over the past decade. These models are effective at learning decision boundaries from large-scale labeled datasets, enabling accurate recognition and judgment of known concepts. However, this intelligence, based on discrimination, has shown limitations in creativity and generalization. It excels at distinguishing known concepts within a single modality but struggles both to generate novel content and to establish semantic associations across multiple modalities. The rise of generative models addresses these limitations by employing large-scale pre-training to capture the underlying data distribution. As a result, mainstream generative models such as Large Language Models (LLMs) [12, 13, 14], Multimodal Large Language Models (MLLMs) [15, 16, 17, 18, 19], Vision-Language Action (VLA) [20, 21], and Diffusion Models [22, 23] are now capable of not only interpreting input features but also actively generating coherent text, realistic images, and even cross-modal content. As the field transitions from discriminative to generative paradigms, AI is evolving from merely recognizing the known to actively creating the new. This shift is redefining the foundations of human-machine collaboration by enabling more flexible, and multimodal generation capabilities.
Figure 1: (a) Continual learning lets humans learn lifelong without forgetting; AI models aim to do the same while retaining old knowledge. (b) Discriminative models focus on perceptual tasks such as classification, detection, and segmentation, among others, whereas generative models additionally demand the capability to produce novel content grounded in perceptual understanding.
The success of generative models like GPT-4o fundamentally relies on large-scale pre-training and human alignment. However, this process is inherently dynamic, as languages, data distributions, and user demands continually evolve. More importantly, it is infeasible to account for all possible future scenarios in advance, which underscores the need for models to possess continual learning capabilities comparable to those of humans, as illustrated in Figure 1 (a). Such capabilities are therefore essential for generative models to move beyond the closed-world assumption [24] and enable deployment in open-ended, real-world scenarios [25, 26, 27]. A central challenge in this context is catastrophic forgetting, where learning new information causes previously acquired knowledge to be overwritten [28, 29, 30]. For instance, fine-tuning general-domain models (e.g., ChatGPT) on specialized domains (e.g., mathematics) may improve in-domain performance at the expense of forgetting prior capabilities. When models are trained sequentially on multiple downstream tasks, each new update compounds this effect, leading to a progressive degradation in knowledge retention. A straightforward solution is to retain data from previous tasks and perform joint training with new data. However, this approach is often impractical due to significant storage costs [31] and data privacy concerns [32]. In contrast, both children and adults rarely experience catastrophic forgetting when acquiring new knowledge [33, 34], highlighting a fundamental gap between artificial and biological learning systems. To bridge this gap, the study of continual learning [35, 36, 37, 38, 39] has gained increasing attention, aiming to equip models with the ability to acquire knowledge sequentially while preserving previously learned skills.
Compared to traditional discriminative models that focused on perceptual tasks, generative models in continual learning settings are expected to go a step further, as shown in Figure 1 (b), by generating new content based on understanding and recognition, owing to the diversity of modeling objectives and task formats. On the one hand, discriminative methods focus on modeling conditional probabilities (i.e. $P(y|x)$), where catastrophic forgetting primarily manifests as a drift in the decision boundary within the classification space [40, 41, 42, 43, 44]. In contrast, generative models aim to learn joint probability distributions (i.e. $P(x,y)$) or perform autoregressive generation, which requires continual learning to maintain both semantic coherence and knowledge completeness within an open-ended generative space. To achieve this, it is essential to refine the generative reasoning process alongside the underlying factual knowledge representation, thereby mitigating hallucinations [45, 46] and semantic inconsistencies [47]. On the other hand, generative models encounter highly diverse tasks in the continual learning process, with differences arising from factors such as new modality inputs, novel domain knowledge, and other evolving task characteristics. This variability increases the risk of catastrophic forgetting and places greater demands on multi-task knowledge integration. In contrast, discriminative models typically focus on a narrow range of tasks, such as classification, detection, or segmentation, which makes continual learning comparatively less challenging. Figure 2 demonstrates the continual learning scenarios across four representative generative models. Despite their impressive capabilities, enabling these models to learn continuously remains an open and pressing research challenge.
Figure 2: Illustrations of continual learning in generative models: Adapting to real-world demands by learning new instructions, knowledge, modalities, actions, and creations without catastrophic forgetting.
Existing continual learning methods have been extensively developed for discriminative models [48, 49, 50, 51, 52]. In contrast, research on continual learning in generative models has generally focused on specific architectures, such as Large Language Models [53, 54, 55, 56] or Multimodal Large Language Models [57, 58, 59]. Although these works provide systematic overviews of continual learning techniques within their respective domains, they often lack broader discussions and cross-model connections. Motivated by this gap, and considering the growing interest in generative and large-scale unified models within the AI community, we believe that an up-to-date and comprehensive survey will offer a holistic perspective to guide future work. In this paper, we present the first comprehensive review of continual learning in mainstream generative models, systematically analyzing key approaches and challenges across Large Language Models, Multimodal Large Language Models, Vision-Language Action Models, and Diffusion Models, as illustrated in Figure 2. In addition, we provide a unified perspective by identifying shared methodological dimensions of these generative models in the context of continual learning.
The paper is organized as follows: In Section 2, we present the setup of continual learning, including its basic formulation, evaluation metrics, and taxonomy. In Section 3 to Section 6, we provide an in-depth discussion of continual learning in Large Language Models (Section 3), Multimodal Large Language Models (Section 4), Vision-Language Action Models (Section 5), and Diffusion Models (Section 6). Each domain is discussed in terms of the training strategies for continual learning tasks, common benchmarks, model architectures, as well as the motivations and implementations of existing methods. In Section 7, we further explore future research avenues and emerging trends in continual learning methods based on generative models. Finally, Section 8 concludes the paper and introduces a continuously updated GitHub repository for tracking recent progress in continual learning for generative models.
2 Overview of Continual Learning
[…]
(§ 3), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=3.0cm [Architecture-based
(§ 3.2), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.7cm [TPEM [60], Zheng et al [61], TGL [62], LOIRE [63], AdapterCL [64], SAPT [65], MoRAL [66], I-LoRA [67], SLIM [68],
P-Prompts [69], Q-Tuning [70], WAVE-CRE [71] , color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
(§ 3.3), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.7cm [O-LoRA [72], DAS [73], ARPER [74], SEEKR [75], TaSL [76],
DYNAINST [77], DaCL [78], FVG [79], color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.5cm] ] [Replay-based
(§ 3.4), color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=2.7cm [CT0 [80], Huang et al. [81], Anh et al. [82], InsCL [83], HESIT [84],
Sun et al. [85], D-CPT Law [86], LAMOL [87], HMI-LAMOL [88], PCLL [89], DCL [90], SSR [91], color=lightcoral!100, fill=lightcoral!15, very thick, text=black, text width=7.5cm] ] ] [Multimodal Large Language Models
(§ 4), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=3.0cm [Architecture-based
(§ 4.2), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [Eproj [92], TAM-CL [93], MoELoRA [94], Continual LLaVA [95], HiDe-LLaVA [96], LLaCA [97], SMoLoRA [98], CL-MoE [99],
LLaVA-CMoE [100], DISCO [101], TRIPLET [102], ModalPrompt [103],
Fwd-Prompt [104], ColPro [105], CluMo [106], MR-LoRA [107], color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
(§ 4.3), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [SCD [108], CS-VQLA [109], Chen et al. [110], CrossSDC [111],
MAFED [112], MoInCL [113], QUAD [114], Model Tailor [115],
SPIDER [116], SEFE [117], color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.5cm] ] [Replay-based
(§ LABEL:MLLM:rep), color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=2.7cm [VQACL [118], SGP [119], ProtoGroup [120], Lin et al. [121],
Adapt-$\infty$ [122], OASIS [123], color=harvestgold!100, fill=harvestgold!15, very thick, text=black, text width=7.5cm] ] ] [Vision-Language Action Models
(§ 5), color=cyan!100, fill=cyan!15, very thick, text=black, text width=3.0cm [Architecture-based
(§ 5.2), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.7cm [LIBERO [124], Lotus [125], QueST [126], LEGION [127], color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
(§ 5.3), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.7cm [CAMA [128], M2Distill [129], color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.5cm] ] [Replay-based
(§ 5.4), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.7cm [iManip [130], RWLA [131], color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.5cm] ] ] [Diffusion Models
(§ 6), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=3.0cm [Architecture-based
(§ 6.2), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.7cm [C-LoRA [132], color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.5cm] ] [Regularization-based
(§ 6.3), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.7cm [MuseumMaker [133], LFS-Diffusion [134], CIDM [135], ConceptGuard [136], Jha et al [137], color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.5cm] ] [Replay-based
(§ 6.4), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.7cm [$\text{L}^{2}\text{DM}$ [138], color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.5cm] ] ] ]
Figure 3: Taxonomy of Continual Learning in Generative Models.
2.1 Basic Formulation
Continual learning tackles the challenge of acquiring knowledge from dynamically evolving data distributions, where training samples from distinct tasks arrive sequentially. A continual learning model is required to assimilate new knowledge while preserving previously learned information, typically under constraints that limit or prevent access to past data, while maintaining strong performance across all tasks. Formally, suppose there are $T$ continual learning tasks in total. The training samples for task $t$ can be represented as $\mathcal{D}_{t}=\{(x,y)\}$, where $x$ is the input data and $y$ is the corresponding labels. Notably, the types of training data required vary across different generative models. For example, in MLLM, $(x,y)$ typically consists of a triplet comprising input images, user instructions, and target answers $(\textbf{X}_{v},\textbf{X}_{ins},\textbf{X}_{a})$. In VLA models, it can take the form of state-action pairs $(\mathcal{S},\mathcal{A})$, where the state $\mathcal{S}$ includes an image observation and a text instruction. During the training of task $t$, only the data for the current task is accessible. The training objective of each continual learning task can be formalized as follows:
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))\right],$ (1)
where $f_{\boldsymbol{\theta}}$ represents the generative model, and $\ell(\cdot,\cdot)$ is the task-specific loss function. The specific formulation for each generative model is detailed at the beginning of its corresponding section.
Upon completion of training for the $t$-th task, the model is evaluated on the test sets of all previously seen tasks, without having access to any task identity information for the current input. An ideal continual learning framework for generative models should enable comprehensive learning of the current task, preserve previously acquired knowledge, and demonstrate strong generalization to future unseen tasks.
2.2 Evaluation Metrics
The evaluation of a model’s continual learning ability typically involves three aspects: overall performance on all learned tasks, the extent of forgetting on previously learned tasks, and the generalization capability to unseen tasks. For clarity, we denote $a_{t,k}$ as the performance on the $k$-th task after the model has been sequentially trained from task 1 to task $t$ [49].
Overall performance typically includes Last Accuracy (Last) and Average Accuracy (Avg). These two metrics are calculated after training on the $t$-th task as follows:
$\text{Avg}_{t}=\frac{1}{t}\sum_{j=1}^{t}\text{Last}_{j},~{}\text{Last}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (2)
where Last reflects the model’s average performance on the current task and Avg further synthesizes the performance across previous tasks. For both metrics, higher values correspond to superior continual learning performance.
Forgetting evaluation typically includes two key metrics: the forgetting measure (FM) and backward transfer (BWT). The former first computes the forgetting degree for each seen task, derived from the difference between the historically highest accuracy and the current accuracy:
$f_{k,t}=\max_{i\in\{1,\cdots,t-1\}}(a_{i,k}-a_{t,k}),\forall k<t.$ (3)
FM is calculated as the average forgetting degree across all previous tasks at task $t$:
$\text{FM}_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}f_{i,t}.$ (4)
BWT quantifies the average impact of learning the $t$-th task on performance across all previously learned tasks:
$\text{BWT}_{t}=\frac{1}{t-1}\sum_{k=1}^{t-1}(a_{t,k}-a_{k,k}).$ (5)
The forgetting measure (FM) quantifies catastrophic forgetting, where lower values indicate better retention of previous knowledge. BWT evaluates knowledge transfer between tasks, with positive values (i.e. BWT > 0) reflecting beneficial forward transfer to prior tasks and negative values (i.e. BWT < 0) indicating interference.
Generalization capability can be evaluated by zero-shot transfer (ZT), which is designed to quantify how well zero-shot transferability is maintained throughout the training process. Formally, for a given task $k$, it first calculates the average performance $A_{k}$ of all models trained on preceding tasks when evaluated on task $k$:
$A_{k}=\frac{1}{k-1}\sum_{i=1}^{k-1}a_{i,k},\forall k>1.$ (6)
ZT is the average of $A_{k}$ after training on the final task:
$\text{ZT}=\frac{1}{T-1}\sum_{k=2}^{T}A_{k}.$ (7)
A higher value of ZT demonstrates superior preservation of zero-shot generalization capability for unseen tasks throughout the continual learning process.
It is worth noting that the calculation of $a_{t,k}$ depends on the specific task type. For example, Bilingual Evaluation Understudy (BLEU) [139] for text generation task, Intersection-over-Union (IoU) [140] for visual grounding task, Fréchet Inception Distance (FID) [141] for text-to-image generation task, and so on. There are several other evaluation criteria available, and we refer readers to their original papers for detailed specifications.
2.3 Taxonomy
Inspired by the collaborative mechanisms of rapid learning in the hippocampus and long-term integration in the neocortex [142, 143, 144, 145, 146], continual learning methods can be systematically categorized into the following three approaches.
Architecture-based methods mimic the brain’s modular organization [147, 148, 149, 150] through dynamic network expansion or modular design to isolate task-specific knowledge. These methods typically encapsulate such knowledge within distinct network modules. As model sizes grow, existing approaches increasingly adopt Parameter-Efficient Fine-Tuning (PEFT) techniques (e.g., LoRA [151] and Prompt tuning [152]) to store and adapt knowledge. Formally, let $\theta_{old}$ denote the model parameters that remain frozen during training on the current task. These typically include the backbone network and the extended sub-network associated with previous tasks. The parameters specific to the current task are denoted by $\theta_{new}$, which are usually the only trainable components. At this point, Eq (1) is rewritten as follows:
$\mathop{\arg\min}_{\theta_{new}}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta_{old}\bigcup\theta_{new}}(x),y))\right].$ (8)
Since the generative model inherently possesses strong generalization capabilities, these sub-modules with relatively few parameters can effectively retain task-specific knowledge. During inference, one of the key challenges for this class of methods is to effectively select the appropriate sub-module based on the input sample.
Regularization-based methods emulate neocortical synaptic stability and metaplasticity by constraining critical parameters to preserve learned representations, mirroring biological selective consolidation [153, 154, 155, 156, 157]. To achieve this goal, generative models often employ regularization in either the parameter space or the feature space, with the former formalized as follows:
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))+\lambda\cdot\Omega(\theta,\theta_{old})\right],$ (9)
where $\lambda$ is a pre-defined hyperparameter, and $\Omega(\cdot,\cdot)$ is a regularization term (e.g., L2 norm) enforcing constraints between the current parameter $\theta$ and the old task parameter $\theta_{old}$.
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))+\lambda\cdot\mathcal{R}(\phi,\phi_{old})\right],$ (10)
where $\phi=f_{\theta}(x)$ denotes the output features or logits extracted by the model, and $\mathcal{R}(\cdot,\cdot)$ is a feature regularization term (e.g., knowledge distillation [158]) used to preserve the representations learned from previous tasks. The central challenge of such methods is to devise regularization terms that effectively retain knowledge from prior tasks with minimal impact on the acquisition of new knowledge.
Replay-based methods replicate hippocampal memory replay mechanisms [159, 160, 161, 162] by storing raw data or generating synthetic samples to revisit past experiences during new task training, effectively mitigating catastrophic forgetting. In practice, this approach typically involves maintaining an auxiliary memory buffer $\mathcal{M}$ that stores a subset of data from previous tasks, which is then used in joint training with current task data:
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim(\mathcal{D}_{t}\bigcup\mathcal{M})}\left[\ell(f_{\theta}(x),y))\right].$ (11)
For generative models, a key challenge in replay-based continual learning methods stems from the limited capacity of memory buffers. Given this constraint, selecting a small yet representative subset of past samples for rehearsal becomes a critical issue. Furthermore, considering practical concerns such as data privacy and storage overhead, several approaches opt to retain intermediate representations (e.g., features or hidden states) of previous tasks rather than storing raw samples, thereby mitigating the risks and costs associated with direct data retention.
Accordingly, we categorize existing generative model-based continual learning approaches according to the taxonomy outlined above, as illustrated in Figure 3. Notably, reflecting the brain’s distributed processing of complex information via interconnected neural units, these methods often consist of multiple interdependent modules. For analytical clarity, each method is primarily classified based on its main functional component, while additional design dimensions are introduced in detail within the corresponding descriptive sections.
3 Continual Learning in Large Language Models
In recent years, Large Language Models (LLMs) [12, 163, 164] have demonstrated outstanding natural language understanding and generation capabilities by pre-training on massive general-domain text corpora. To sustain high performance in real-world applications, LLMs require the capability to integrate new information and retain prior knowledge as data, tasks, and user preferences evolve. Unlike traditional language models, LLMs treat every task as a generative sequence prediction problem through autoregressive next-token prediction. Their massive parameterization enables emergent capabilities including few-shot learning [165] and complex multi-step reasoning [166]. Consequently, many conventional continual learning methods designed for smaller models struggle to adapt effectively to LLMs.
In the following section, we first outline the setup for an LLM continual learning task (Section 3.1), including its training objectives, commonly used model backbones, and widely adopted benchmarks. Next, we classify the existing research into three strategies: architecture-based (Section 3.2), regularization-based (Section 3.3), and replay-based (Section 3.4) approaches.
3.1 Problem Setup
3.1.1 Training objectives
In the instruction tuning paradigm for a LLM $f_{\theta}$, the training dataset $\mathcal{D}=\{(x_{j},y_{j})\}_{j=1}^{N}$ typically consists of input texts $x$ (e.g.,, questions, instructions, or context) paired with target output texts $y$ (e.g.,, answers, summaries, or responses), where $N$ denotes the total number of samples. Given an input $x$ and its corresponding output $y$ of length $L$, the model generates $y$ autoregressively:
$p(y\mid x)=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(y_{i}\mid x,y_{<i}),$ (12)
where $\theta$ denotes the trainable parameters during fine-tuning, $y_{i}$ is the $i$-th token in the output sequence, and $y_{<i}$ represents all previously generated tokens before position $i$. The objective of instruction tuning is to minimize the overall loss on the dataset $\mathcal{D}$:
$\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y})\sim\mathcal{D}}\left[\ell(f_{\boldsymbol{\theta}}(\boldsymbol{x}),\boldsymbol{y})\right],$ (13)
where $x$ and $y$ denote the input and target texts, respectively, and $\ell(\cdot,\cdot)$ is the loss function, typically the token-level cross-entropy loss, which measures the discrepancy between the predicted output and the ground-truth target.
Figure 4: Overview of continual learning approaches in Large Language Models.
In a continual learning scenario, the training data for the model arrives as a continuous stream $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{T}\}$, comprising a sequence of $T$ tasks. The objective of continual learning for LLM is to incrementally train on these $T$ datasets, aiming to maximize performance on the current task while preserving knowledge from previous tasks. Formally, at task $t$, the optimization objective in Eq. (13) can be reformulated as:
$\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y})\sim\mathcal{D}}\left[\ell(f_{\boldsymbol{\theta}_{t}}(\boldsymbol{x}),\boldsymbol{y})\right]+\sum_{i=1}^{t-1}\epsilon_{i},$ (14)
[…]
$\epsilon_{i}\geq 0;\forall i\in[1,\cdots,t-1],$
where $\epsilon_{i}$ is a slack variable that permits a slight increase in loss on previous tasks, allowing controlled forgetting while prioritizing learning on the current task.
3.1.2 Backbone
In natural language generation tasks, mainstream LLMs generally fall into two architectural categories: encoder-decoder architectures and decoder-only architectures. According to recent literature, we summarize the backbone architectures commonly adopted in LLM’s continual learning task as follows:
  • BERT [167] employs a bidirectional encoder-only architecture based on the Transformer framework. Unlike autoregressive models, BERT leverages masked language modeling and next-sentence prediction as its pretraining objectives, enabling it to learn contextualized word representations from both left and right contexts simultaneously. Each encoder layer comprises multi-head self-attention, feedforward networks, layer normalization, and residual connections.
  • T5 [168] adopts a standard encoder-decoder framework, where both the encoder and decoder are built upon the Transformer architecture. T5 reformulates all NLP tasks into a unified text-to-text format. Each layer in T5 consists of conventional multi-head attention and feedforward sublayers, with configurable embedding dimensions and attention heads depending on the model scale.
  • GPT [169] series exemplifies a unidirectional decoder-only architecture, characterized by left-to-right autoregressive generation. GPT is composed of stacked decoder blocks, each containing masked multi-head attention, feedforward layers, and residual connections. The design is streamlined and restricts visibility to a leftward context, making it well-suited for open-ended generation tasks. In most research settings, the GPT-2 model is used.
  • LLaMA [13] series retains the decoder-only design while introducing several architectural optimizations, including Rotary Positional Embeddings (RoPE), the SwiGLU activation function, and RMSNorm. These improvements bolster training stability, accelerate inference, and preserve performance on tasks requiring extended context lengths.
In addition to the mainstream backbone networks mentioned above, some approaches use other architectures as well, and we refer readers to their original papers.
3.1.3 Benchmark
Several benchmark frameworks exist to evaluate the continual learning performance of LLMs. In this part, we focus on some of the most representative benchmarks to provide a detailed introduction.
  • Short-sequence CL  [170] comprises five large-scale text classification datasets, including AG News (news classification, 4 classes), Amazon Reviews (sentiment analysis, 5 classes), Yelp Reviews (sentiment analysis, 5 classes), DBpedia (Wikipedia text classification, 14 classes), and Yahoo Answers (Q&A classification, 10 classes).
  • Long-sequence CL  [69] builds on the five datasets above by adding four tasks from the GLUE benchmark (MNLI, QQP, RTE, SST-2), five tasks from SuperGLUE (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB dataset, forming a longer continual learning evaluation sequence.
  • SuperNI [171] is a large-scale benchmark comprising 1,616 NLP tasks paired with natural language instructions, spanning 76 task types and 55 languages. Each task is accompanied by a clear task definition and illustrative examples demonstrating the mapping from input to the desired output. A representative continual learning workflow is often simulated using a diverse subset of tasks selected across domains and types from this benchmark.
  • TRACE [172] includes eight challenging datasets spanning domain-specific knowledge, multilingual ability, code generation, and mathematical reasoning. All datasets are uniformly formatted to allow automated evaluation. To measure forgetting in aligned LLMs during continual learning, TRACE introduces three core metrics: (1) General Ability Delta: tracks changes in performance on general tasks such as factual QA, reasoning, multilingual QA, commonsense reasoning, and reading comprehension; (2) Instruction Following Delta: measures shifts in instruction adherence on Self-instruct and LIMA prompt sets; (3) Safety Delta: evaluates changes in response safety using the CoNa dataset.
TABLE I: Summary of continual learning methods for Large Language Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Setting: CToD refers to Continual Task-oriented Dialogue, CIT to Continual Instruction Tuning, CPT to Continual Pre-Training, and CRE to Continual Relation Extraction.
Methods Taxonomy Setting Backbone / Basic models Implement Code
Arch Reg Rep
TPEM [60] CToD GLMP [173] Backbone Link
Zheng et al. [61] CIT LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, Pythia-410M [174] Backbone Link
TGL [62] CPT GPT-2 base, GPT-2 large, GPT-Neo Backbone -
LOIRE [63] CPT GPT-2 small, GPT-2 large, BERT-base Backbone -
AdapterCL [64] CToD GPT-2 PEFT Link
SAPT [65] CIT T5-large, T5-XL, T5-XXL, LLaMA-2-7B, LLaMA-2-13B PEFT Link
MoRAL [66] CPT TinyLlama-1.1B [175], Phi-2-2.7B [176], LLaMA-2-7B PEFT -
I-LoRA [67] CIT LLaMA-2-7B PEFT Link
SLIM [68] CIT OpenChat-8B [177] PEFT -
P-Prompts [69] CIT BERT-base, T5-large PEFT Link
Q-Tuning [70] CIT BERT-base, T5-large PEFT -
WAVE-CRE [71] CRE BERT-base PEFT Link
O-LoRA [72] CIT T5-large, LLaMA-7B PEFT Link
DAS [73] CPT RoBERTa-base [178] Backbone Link
ARPER [74] CToD SCLSTM [179] Backbone Link
SEEKR [75] CIT LLaMA-2-7B-Chat, Vicuna-7B-v1.5 [180], Vicuna-13B-v1.5 Backbone Link
TaSL [76] CToD T5-small, T5-base, T5-large, LLaMA-7B Backbone & PEFT Link
DYNAINST [77] CIT BART-base [181] Backbone -
DaCL [78] CRE BERT-base Backbone Link
FVG [79] CIT LLaMA-2-7B-Chat, LLaMA-2-13B-Chat, LLaMA-3-8B-Chat, Mistral-7B-instruct-v2 [182] PEFT -
CT0 [80] CIT T5-small, T5-3B Backbone Link
Huang et al. [81] CIT OPT-125M [183], OPT-350M, OPT-1.3B, OPT-2.7B, OPT-6.7B, OPT-13B Backbone -
Anh et al. [82] CRE BERT-base, LLM2Vec [184] Backbone -
InsCL [83] CIT LLaMA-7B Backbone Link
HESIT [84] CToD GPT-2 Backbone -
Sun et al. [85] CIT Qwen2-0.5B [164], Qwen2-1.5B, LLaMA-2-7B, LLaMA-2-13B, Mistral-7B [182] Backbone Link
D-CPT Law [86] CPT Qwen1.5-0.5B [185], Qwen1.5-1.8B, Qwen1.5-4B Backbone -
LAMOL [87] CIT GPT-2 small Backbone Link
HMI-LAMOL [88] CIT GPT-2 small, BERT-base Backbone Link
PCLL [89] CToD GPT-2 small PEFT Link
DCL [90] CToD GPT-2 small PEFT -
SSR [91] CIT LLaMA-2-7B, LLaMA-2-7B-Chat, Alpaca-7B [186] Backbone Link
3.2 Architecture-based Approach
3.2.1 Backbone Expansion
Backbone expansion methods dynamically adapt network architectures to integrate new information by modifying parameter sparsity, expanding model capacity, or introducing new learning modules, without requiring full retraining. As shown in Figure 4 (a), these methods typically restructure internal components to balance the retention of previous knowledge with the acquisition of new capabilities. For instance, TPEM [60] proposes a continual learning framework for task-oriented dialogue systems based on iterative pruning, expansion, and masking. It begins by pruning redundant weights and retraining the retained parameters to achieve sparsity without sacrificing performance. To support new tasks, TPEM dynamically expands hidden dimensions based on pruning ratios and data volume. By analyzing early-stage optimization dynamics, Zheng et al. [61] reveal that performance drops often stem from disrupted alignment rather than erased knowledge. A layer-freezing strategy is proposed to stabilize low-level representations, yielding consistent improvements across multiple CL benchmarks. TGL [62] proposes a gradient-probing approach to localize time-sensitive entity knowledge in LLMs using salient span prediction tasks. Compared to generic continual pretraining or knowledge editing techniques, this method emphasizes the locality of knowledge in model parameters, demonstrating that selective layer updates can enable more efficient knowledge adaptation and mitigate catastrophic forgetting. LOIRE [63] introduces a plug-in layer expansion strategy with residual connections to ensure strict function preservation during depth-wise scaling. It further optimizes model growth across hidden dimensions, feed-forward layers, and attention heads. To address catastrophic forgetting, LOIRE employs an iterative distillation strategy, where intermediate models alternately act as teachers and students, enabling stable knowledge retention across evolving data distributions.
3.2.2 LoRA Expansion
To enable effective knowledge acquisition while preserving the backbone network’s original capabilities, parameter-efficient fine-tuning methods like LoRA and Adapter [187] have been widely adopted. These approaches leverage task-specific low-rank updates to the weight matrices, allowing the model to adapt to new tasks while maintaining stable performance on previously learned ones, thereby mitigating catastrophic interference. For example, AdapterCL [64] utilizes residual adapters to separate task-specific knowledge and adopts a perplexity-based inference strategy for adapter selection. SAPT [65] introduces a framework that aligns the learning and selection of PET blocks via a Shared Attentive Learning and Selection module to tackle catastrophic forgetting and knowledge transfer. It employs an Attentive Reflection Module to leverage generated pseudo samples for recalling shared attention from prior tasks, ensuring robust performance without requiring task identifiers during testing. MoRAL [66] integrates LoRA with mixture-of-experts routing to enable parameter-efficient adaptation and dynamic expert selection. I-LoRA [67] builds a dual-memory replay framework with a fast learner for quick adaptation and a slow learner for long-term knowledge consolidation. SLIM [68] introduces a dynamic mechanism that selectively activates or bypasses LoRA modules based on input characteristics. Specifically, it employs a learnable routing layer and sliding clustering to distinguish in-domain and out-of-domain inputs, directing general tasks through identity mappings and downstream tasks through LoRA adapters.
3.2.3 Prompt Expansion
As an alternative parameter-efficient fine-tuning approach for LLMs, prompt-based methods achieve continual learning by modularizing knowledge into prompts [188, 152], enabling efficient task adaptation and selective updates while reducing forgetting and interference. P-Prompts [69] learn a distinct prompt for each task and progressively concatenate them before feeding into a frozen backbone. It prevents forgetting by keeping task-specific prompts fixed after training while enabling forward transfer through the reuse of previous prompts. Q-Tuning [70] proposes a continual prompt tuning framework that maintains a prompt queue for old tasks and introduces new prompts for incoming tasks. It applies a low-rank adaptive knowledge aggregation mechanism for relevance reweighting, employs a PCA-based eviction rule to maintain constant queue size, and incorporates a globally shared prefix prompt with memory retention regularization to mitigate information loss across tasks. WAVE-CRE [71] proposes a continual relation extraction method that combines task-specific prompt pools with generative consolidation. The approach introduces a prompt pool for each task to capture intra-task variations and leverages a generative model to consolidate prior knowledge within shared parameters, effectively mitigating catastrophic forgetting without explicit data storage.
3.3 Regularization-based Approach
Regularization-based approaches, as illustrated in Figure 4 (b), mitigate catastrophic forgetting by enforcing training-time constraints that preserve critical parameters or feature representations from prior tasks. For example, O-LoRA [72] employs task-specific LoRA modules with orthogonal regularization, where each new task receives dedicated low-rank adaptation parameters. By enforcing orthogonality constraints on the LoRA updates, it effectively minimizes interference between task-specific parameter spaces, thereby substantially mitigating catastrophic forgetting. DAS [73] introduces a continual domain-adaptive pre-training framework for language models, employing a soft-masking mechanism to regulate parameter updates based on the importance of units for general or domain-specific knowledge. It preserves general knowledge through a robustness-based proxy and facilitates cross-domain transfer by contrasting learned knowledge with complete knowledge, thereby mitigating catastrophic forgetting and enabling knowledge integration without requiring domain IDs during fine-tuning. ARPER [74] applies an adaptive elastic weight consolidation regularization [30] to dynamically constrain important parameters based on vocabulary shifts across tasks, assigning stronger penalties to parameters critical for previous knowledge. SEEKR [75] employs knowledge distillation [158] on attention head outputs to preserve critical knowledge, while introducing task sensitivity and forgettability measures to dynamically identify and prioritize the most informative attention heads for efficient knowledge retention. TaSL [76] employs group-wise gradient-based importance scoring to identify task-specific and shared parameters, followed by fine-grained model averaging that selectively integrates knowledge from previous tasks while preserving critical parameters to mitigate forgetting and enhance bidirectional knowledge transfer. DYNAINST [77] is a hybrid method for lifelong in-context instruction learning that combines wide local minima regularization with dynamic instruction replay. Specifically, it introduces a KL-divergence term that encourages the model’s output distribution to remain smooth and less overconfident, thereby improving robustness and generalization to future tasks. NaCL [78] presents a noise-resistant contrastive learning framework for continual relation extraction, employing a rebooted auxiliary model to identify clean samples and performing noise-guided adversarial attacks to align noisy samples with their labels in feature space, enriching contrastive representations and mitigating forgetting. FVG [79] proposes function vector-guided training [189] that stabilizes function vectors through activation consistency and KL-divergence alignment. It reveals that forgetting primarily arises from biased function activations rather than parameter overwriting, positioning function vectors as effective tools for diagnosing forgetting.
3.4 Replay-based Approach
3.4.1 Direct Replay
An intuitive replay-based approach involves retaining selected data samples from previous tasks or their compressed representations. As shown in Figure 4 (c), these stored exemplars are strategically reintroduced during the training of new tasks to prevent catastrophic forgetting. For example, CT0 [80] introduces a rehearsal mechanism by replaying 1% of historical data from each task, which allows the model to continually adapt to new tasks while preserving its previously acquired instruction-following ability. Huang et al. [81] propose a tool-augmented mechanism combined with a small replay buffer, which helps improve the model’s adaptation to new tasks while effectively mitigating the forgetting of old tasks, thus enhancing overall performance across both seen and unseen tasks. Anh et al. [82] combine old and new task samples into complex dual-label texts to mitigate data scarcity. They enhance model robustness through adversarial-like perturbations and employ specialized loss functions to consolidate relation representations while reducing overfitting. The method further improves generalization by selectively applying Sharpness-Aware Minimization [190] to current-task training data.
Moreover, several methods adopt dynamic data replay strategies. InsCL [83] adjusts the replay ratio of historical data dynamically based on task similarity. It also proposes a new metric for measuring instruction informativeness, favoring high-quality samples with richer information for training. HESIT [84] tracks how much each sample changes the model early in training (i.e. the gradient of model parameters with respect to the sample’s weight), then picks and replays the most impactful examples to effectively reduce forgetting of old tasks. Sun et al. [85] demonstrate that performance degradation in continual learning stems from the original prompts’ inability to effectively steer the model’s rationale generation. They introduce the Rationale-Guidance Difficulty metric to measure how hard it is for a prompt to guide reasoning, and they use this metric to replay old data as needed to keep and restore past task performance. There are also works that aim to optimize the allocation ratio between replay data and current task data in continual learning. For example, D-CPT Law [86] establishes a scaling law to optimize the mixture ratio between general and domain-specific corpora. This method extends the Chinchilla scaling law [191] by incorporating mixture ratio as a parameter and introducing a Domain-specific Learnable Coefficient to enable cross-domain generalization in limited-data scenarios.
3.4.2 Generative Replay
Generative replay methods leverage LLMs to synthesize pseudo-samples from previous tasks, enabling memory consolidation without needing to store real historical data. For example, LAMOL [87] introduces a generation token to guide the model in producing pseudo-samples alongside new task learning. This approach enhances the model’s ability to retain old knowledge without incurring extra computational costs. HMI-LAMOL [88] proposes a brain-inspired generative replay mechanism, HMI, which uses BERT to encode training samples into compressed “memory traces”. These traces condition the generation of old task data, effectively alleviating sample imbalance and quality degradation, thus mitigating catastrophic forgetting in language models. Furthermore, PCLL [89] constructs natural language prompts for each task and employs a conditional variational autoencoder to model the task-specific latent distribution, generating high-quality pseudo-samples for replay. DCL [90] replaces the Gaussian distribution used in PCLL with a Dirichlet distribution to model the latent variables, enabling better representation of structural differences across tasks, and introduces Jensen–Shannon knowledge distillation at the logit layer to improve the model’s retention of old knowledge. SSR [91] further introduces a self-synthesizing reproduction framework, which utilizes the context-learning ability of a base LLM to generate synthetic samples. These samples are refined by the latest fine-tuned model, and high-quality ones are selected via clustering algorithms, enabling the retention of prior knowledge without requiring historical data or additional generative models.
3.5 Discussions
This section reviewed the application of continual learning in large language models, focusing on three main approaches: architecture-based, regularization-based, and replay-based methods. Among them, architectural expansion and replay have emerged as dominant strategies due to the task diversity and semantic complexity inherent to LLMs. The former enhances task specialization by introducing lightweight, modular components such as plug-in LoRA adapters, offering scalability and modularity. Replay-based methods, on the other hand, leverage the strong generalization capabilities of LLMs, requiring only a small number of previous samples or features to effectively mitigate forgetting. While regularization has proven effective in traditional networks, its adoption in LLMs remains limited due to challenges in tuning and stability. Nonetheless, the integration of regularization with architectural expansion or replay strategies, such as applying knowledge-preserving constraints during multitask pretraining or dynamically evaluating parameter importance during updates, holds great potential for developing more robust and generalizable continual learning frameworks in LLMs.
4 Continual Learning in Multimodal Large Language Models
Recent Multimodal Large Language Models [192], built upon advanced Large Language Models, have demonstrated impressive visual understanding capabilities by integrating visual encoders, cross-modal projectors, and leveraging large-scale image-text corpus during training. Continual learning in MLLMs seeks to equip them with task-specific abilities across domains such as healthcare [193, 194], finance [195, 196], and autonomous driving [197, 198], without forgetting their previously learned knowledge. In contrast to discriminative models that predict output through classifiers, MLLMs generate responses by next-token prediction in an autoregressive way, making most continual learning methods designed for discriminative models less applicable. Additionally, compared to LLMs, MLLMs need to simultaneously process information from multiple modalities [199], thus posing a greater challenge.
In this section, we begin by introducing the setup for an MLLM continual learning task (Section 4.1), which includes training objectives, common model backbones, and existing benchmarks. Subsequently, the existing literature is categorized into three strategies: architecture-based (Section 4.2), regularization-based (Section 4.3), and replay-based (Section 4.4) approaches.
4.1 Problem Setup
4.1.1 Training objectives
The visual instruction tuning dataset $\mathcal{D}=\{(\textbf{X}_{v}^{j},\textbf{X}_{ins}^{j},\textbf{X}_{a}^{j})_{j=1}^{N}\}$ for fine-tuning a MLLM $f_{\boldsymbol{\theta}}$ typically consists of image input $\textbf{X}_{v}$, user instruction $\textbf{X}_{ins}$ and target answer $\textbf{X}_{a}$, where $N$ represents the total number of training samples. Given an image-instruction pair with a response of length $L$, the MLLM predicts the next token autoregressively based on all preceding tokens:
$p(\textbf{X}_{a}|\textbf{X}_{v},\textbf{X}_{ins})=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(x_{i}|\textbf{X}_{v},\textbf{X}_{ins},\textbf{X}_{a,<i}),$ (15)
where $\theta$ denotes the trainable parameters during fine-tuning, $\textbf{X}_{a,<i}$ are the response tokens from all previous turns before the current prediction token $x_{i}$. In the continual learning framework for MLLMs, the optimization objective and loss function maintain the same formulation as the LLM case in Eq.(13) and Eq.(14), with the distinction that the input-output pairs now involve multimodal data. Here, $\boldsymbol{x}$ and $\boldsymbol{y}$ represent the multimodal inputs $(\textbf{X}_{v},\textbf{X}_{ins})$ and the target answer $\textbf{X}_{a}$, respectively.
4.1.2 Backbone
To handle inputs from multiple modalities, MLLMs typically begin by extracting modality-specific features using dedicated encoders. These features are subsequently concatenated and fed into a pre-trained language model, which generates responses in an autoregressive manner. Based on existing literature, we summarize the backbone architectures commonly used in continual learning methods for MLLMs as follows.
  • VL-T5 [118] employs two distinct visual encoders for different modalities: Faster R-CNN [8] for extracting region features in visual question answering, and 3D ResNeXt-101 [200] for capturing video motion features in video question answering. The language model in VL-T5 comprises 12 encoder and decoder blocks, each containing a multi-head attention layer with 12 heads, and the embedding dimension is set as 768.
  • InstructBLIP [201] utilizes a Query Transformer (Q-Former) to extract visual features from a frozen image encoder (ViT-G/14 [202]). The input to the Q-Former contains a set of $K$ learnable query embeddings, which interact with the image encoder’s output through cross-attention. The Q-Former outputs are then passed through a linear projection before being fed into the frozen LLM. For the LLM, existing methods [92, 104] typically employ FlanT5-XL which is an instruction-tuned model based on the encoder-decoder Transformer T5 [168]. During training, only the projection layer of the current task in the Q-Former and its corresponding key are trainable.
  • LLaVA [15, 18] series has emerged as one of the most prominent MLLMs, distinguished by its impressive visual understanding capabilities and streamlined architecture. It employs Vicuna [203] as LLM and pre-trained CLIP visual encoder ViT-L/14@336 [204] to extract visual embeddings from images. To align visual and textual modalities, LLaVA employs a simple yet effective linear projection layer to project visual features into a unified representation space.
In addition to the MLLM backbones mentioned above, some methods incorporate customized architectural designs tailored to their specific settings (e.g., Qwen-VL series [17, 19], MiniGPT-v2 [16], InternVL [205], etc.), which are detailed in Table II.
Figure 5: Three types of continual learning approaches in Multimodal Large Language Models.
4.1.3 Benchmark
To evaluate and reveal the continual learning capabilities of MLLMs, several benchmarks have been developed. In this part, we focus on a selection of the most representative benchmarks to provide a detailed discussion. For details regarding the benchmarks and datasets employed by each method, we refer readers to their original papers.
  • CLiMB [206] is introduced to investigate continual learning in vision-and-language settings, supporting both multimodal and unimodal tasks. It consists of two phases: (1) Upstream Continual Learning of Multimodal Tasks: The pre-trained multimodal model is trained on a sequence of vision-and-language tasks and evaluated after each task to assess Forgetting of previous tasks and Knowledge Transfer to subsequent tasks. (2) Downstream Transfer to Low-Shot Tasks: After each multimodal task, the model is evaluated for its Low-Shot Transfer capability on both multimodal and unimodal downstream tasks. The experiments in CLiMB demonstrate that existing continual learning strategies struggle to generalize across sequences of multimodal tasks and fail to enable effective low-shot adaptation to downstream multimodal or unimodal tasks.
  • VQACL [118] designs dual-level task sequence using two datasets: VQA2.0 [207] and NExT-QA [208]. The outer level consists of $T$ linguistic-driven tasks, where each task corresponds to learning a specific reasoning skill. For example, for the ’Count’ task, its training data mainly contain the examples that can teach the model how to count, such as ‘How many’ and ‘What number is’. At the inner level, each linguistic-driven task is further composed of a sequence of visual-driven subtasks, with each subtask containing images from a specific object group. Furthermore, VQACL focuses on evaluating the compositionality of the model’s reasoning skills (e.g., counting, color) and visual concepts (e.g., cat, surfboard). Specifically, a visual-driven subtask is randomly removed from each linguistic-driven task during training and used as a new composition for testing.
  • CoIN [94] consists of carefully designed datasets covering 8 tasks, including general image question answering, visual reasoning, knowledge-grounded image question answering, etc. Additionally, CoIN proposes to evaluate MLLMs from two perspectives: (1) Truth Alignment: The ability to generate the correct result in the desired format, adhering to task instructions, is a fundamental requirement for instruction tuning. (2) Reasoning Capability. The performance of MLLMs depends not only on their ability to follow instructions but also on the knowledge they retain and integrate over time.
  • MLLM-CL [107] introduces two settings named domain continual learning (DCL) and ability continual learning (ACL), providing a comprehensive and realistic evaluation for continual learning of MLLMs. Specifically, DCL focuses on learning domain-specific knowledge continually on several mainstream domains (remote sensing, medical, autonomous driving, science, and finance), and evaluated on independently and identically distributed (IID) test sets. Differently, ACL aims to incorporate fundamental abilities (OCR, math & logic, visual perception and, GUI agent), evaluating non-IID scenarios.
4.2 Architecture-based Approach
4.2.1 Backbone Expansion
As shown in Figure 5 (a), backbone expansion methods dynamically extend model architectures through modular components, enabling adaptive knowledge acquisition for diverse tasks. Eproj [92] establishes the first continual instruction tuning benchmark for MLLM. To address catastrophic forgetting, the framework dynamically selects optimal Q-Former projection layers through multimodal similarity scoring between input samples and preserved training features, while integrating adaptive regularization constraints with a strategic old-sample replay mechanism to maintain stable model performance. TAM-CL [93] proposes a novel transformer-based continual learning framework that dynamically expands model capacity through task-specific attention layers with learnable tokens. The approach simultaneously preserves acquired knowledge through intermediate teacher-student distillation while employing a compact memory buffer for selective experience replay, ensuring robust performance across sequential tasks.
TABLE II: Summary of continual learning methods for Multimodal Large Language Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Benchmark: Composite benchmark refer to method that combine multiple datasets without a unified name. Tag: I = Image, T = Text, V = Video, A = Audio.
Methods Taxonomy Benchmark Backbone / Basic Models Modality Implement Code
Arch Reg Rep
Eproj [92] Composite InstrutBLIP, BILP2 [209] I T Backbone -
TAM-CL [93] Composite ViLT [210] I T Backbone Link
MoELoRA [94] CoIN LLaVA-1.5-7B, LLaVA-1.5-13B, Qwen-VL-Chat, MiniGPT-v2 I T PEFT Link
Continual LLaVA [95] COAST LLaVA-1.5-7B I T PEFT Link
HiDe-LLaVA [96] CoIN, UCIT LLaVA-1.5-7B, InternVL-Chat-7B I T PEFT Link
LLaCA [97] CoIN LLaVA-1.5-7B, LLaVA-1.5-13B, Qwen-VL-Chat I T PEFT Link
SMoLoRA [98] CVIT LLaVA-1.5-7B I T PEFT -
Cl-MoE [99] VQACL VL-T5, LLaVA-1.5-7B I T PEFT Link
LLaVA-CMoE [100] CoIN LLaVA-1.5-7B I T PEFT -
DISCO [101] FCIT LLaVA-1.5-7B I T PEFT -
TRIPLET [102] Composite ALBEF [211], FLAVA [212] I T PEFT -
ModalPrompt [103] CoIN LLaVA-1.5-7B I T PEFT Link
Fwd-Prompt [104] Composite InstructBLIP, BLIP2 I T PEFT -
ColPro [105] Composite LLaMA-7B, CLIP ViT-L/14 V T PEFT Link
CluMo [106] CLOVE ALBEF I T PEFT Link
MR-LoRA [107] MLLM-CL LLaVA-1.5-7B I T PEFT Link
SCD [108] MDL-VQA ViLT I T Backbone -
CS-VQLA [109] Composite ResNet [3], VisualBERT [213] I T Backbone Link
Chen et al [110] Composite InstructBLIP I T Backbone -
CrossSDC [111] Composite iQuery [214] V A Backbone Link
MAFED [112] Composite UNITER [215], ViLT, VL-Pythia I T Backbone Link
[…]
QUAD [114] VQACL VL-T5 I T Backbone Link
Model Tailor [115] Composite InstructBLIP, LLaVA-1.5-7B I T Backbone & PEFT Link
SPIDER [116] Composite VILA-1.5-3B, LLaVA-1.5-7B I T Backbone -
SEFE [117] CoIN-ASD LLaVA-1.5-7B I T PEFT Link
VQACL [118] VQACL VL-T5 I T Backbone Link
SGP [119] CLOVE UniVQA I T Backbone Link
ProtoGroup [120] VQACL VL-T5 I T Backbone -
Lin et al [121] Composite EM-VLM4AD [217] I T Backbone -
Adapt-$\infty$ [122] Composite LLaVA-1.5-7B I T PEFT Link
OASIS [123] MICVIT LLaVA-1.5-7B, Qwen-VL-2.5-7B I T PEFT -
4.2.2 LoRA Expansion
LoRA expansion approaches enhance MLLMs through lightweight low-rank adaptations, where task-specific adapter modules are added to a fixed backbone network, optimally balancing learning plasticity with memory stability for efficient continual learning with minimal forgetting. MoELoRA [94] implements LoRA modules as experts in a mixture-of-experts architecture, employing a learnable router for dynamic expert selection based on task inputs. The approach is evaluated on CoIN, a novel continual instruction tuning benchmark. Continual LLaVA [95] introduces the COAST benchmark, systematically evaluating continual instruction tuning across three dimensions: domain shifts, capability expansion, and dataset evolution. The framework tackles catastrophic forgetting through a novel dual-embedding mechanism, combining instruction-aware intrinsic embeddings with cross-task contextual embeddings, while maintaining parameter efficiency through selective LoRA modules. HiDe-LLaVA [96] proposes a hierarchical decoupling framework for continual instruction tuning in MLLMs. The method analyzes CKA similarity variations across model layers to decompose the architecture into task-specific expansion and task-general fusion, where dual-modality prototype matching guides adaptive LoRA selection while parameter-efficient merging preserves shared knowledge. The work additionally constructs a UCIT benchmark by filtering zero-shot underperforming datasets to eliminate information leakage, enabling a more reliable evaluation of catastrophic forgetting mitigation. MR-LoRA [107] employs explicit architectural decoupling for plasticity preservation while preventing catastrophic interference during continual learning. Furthermore, a multimodal routing mechanism is incorporated, which leverages the MLLM’s intrinsic cross-modal representations to dynamically route inputs to specialized parameters, acting as an explicit expert selector in complex multimodal scenarios. LLaCA [97] introduces a continual instruction tuning framework that dynamically balances stability and plasticity via gradient-guided EMA weight adaptation. By leveraging Taylor expansion for optimal weight determination and semantic-aware instruction grouping, the method achieves strong anti-forgetting while preserving new-task performance. SMoLoRA [98] introduces a separable mixture of low-rank adaptation framework to address dual catastrophic forgetting in continual visual instruction tuning. This method employs distinct routing modules for visual understanding and instruction following, dynamically selecting specialized LoRA blocks through adaptive fusion to preserve both capabilities while learning new tasks. Additionally, this work proposes a CVIT benchmark that evaluates model generalization to unseen tasks and robustness under diverse instructions. CL-MoE [99] introduces a Dual-Router MoE that combines instance-level and task-level routing to dynamically select optimal experts, coupled with a Momentum MoE mechanism that employs differential parameter updating strategies for task-shared and task-specific experts to balance knowledge retention and acquisition. LLaVA-CMoE [100] proposes Probe-Guided Knowledge Extension to dynamically expand experts at necessary layers through activation frequency analysis, coupled with a Probabilistic Task Locator that employs hierarchical routing and VAE-based task distribution modeling to prevent interference between new and existing tasks while preserving zero-shot capabilities. FCIT [101] proposes DISCO, a framework for Federated Continual Instruction Tuning in MLLMs, featuring dynamic knowledge organization to manage task-specific subspaces via identity token matching and subspace selective activation to filter irrelevant outputs during inference. It establishes the first FCIT benchmark, addressing real-world challenges of data heterogeneity and continual learning.
4.2.3 Prompt Expansion
Prompt expansions are a standard technique used to improve the continual learning capability of MLLMs. For example, TRIPLET [102] presents the first multimodal prompt learning framework for continual visual question answering, introducing decoupled prompts across modalities, transformer layers, and tasks, along with query-matching and cross-modal interaction strategies to model complex interactions while mitigating catastrophic forgetting without requiring rehearsal buffers.
ModalPrompt [103] introduces a dual-modality guided prompt learning framework for continual learning of MLLM. It constructs task-specific prototype prompts to retain prior knowledge without data replay and employs a dual-modality (image-text) guided prompt selection mechanism to dynamically match relevant prompts during inference. Fwd-Prompt [104] leverages prompt tuning with gradient projection to achieve both anti-forgetting and positive forward transfer. It addresses task interference through orthogonal subspace allocation, projecting prompt gradients into residual spaces to preserve old knowledge while reusing pre-trained knowledge in core spaces to enhance future task adaptation. ColPro [105] introduces collaborative prompting, which integrates task-specific question constraint prompting, knowledge acquisition prompting, and visual temporal awareness prompting into a pre-trained LLM. It utilizes a complementary learning principle with General and Expert prompts across LLM layers, employing negative guiding for question awareness, autoregressive temporal dynamics for video understanding, and multimodal distillation to enhance task-specific question answering while mitigating catastrophic forgetting. CluMo [106] proposes a cluster-based modality fusion prompt method, which introduces task-specific visual and textual prompt keys trained via K-means clustering in the first stage and uses their combinations to select optimal prompts from a shared pool during the second stage. It employs a two-stage training strategy with knowledge distillation to enhance generalization and mitigate catastrophic forgetting while maintaining a frozen backbone model.
4.3 Regularization-based Approach
4.3.1 Feature-Level Regularization
To overcome catastrophic forgetting, the idea of Knowledge Distillation (KD) has also demonstrated strong potential in continual learning approaches for MLLM. One line of work uses current task data directly for distillation, aiming to retain previous knowledge without relying on past data. For example, SCD [108] distills both instance-relevant and domain-relevant knowledge by analyzing the teacher’s responses to original and counterfactual samples. It employs metric learning to guide the student model in avoiding harmful domain-specific features while effectively acquiring transferable knowledge from new domains. CS-VQLA [109] introduces rigidity-plasticity-aware distillation and self-calibrated heterogeneous distillation to strike a balance between knowledge retention and adaptability, thereby enhancing the preservation of prior knowledge in continual learning for robotic surgery scenarios. Chen et al. [110] introduce an LLM-assisted distillation loss that utilizes the strong generalization ability of LLMs to mitigate the effects of domain shifts and data imbalances. CrossSDC [111] preserves cross-modal associations by maintaining semantic similarity across modalities over time. It can be integrated into various sound separation frameworks to effectively mitigate catastrophic forgetting in continual audio-visual sound separation.
Another line of work leverages distillation with data from previous tasks or generated pseudo-data. Specifically, MAFED [112] mitigates forgetting by mixing a small portion of old samples into the new task data for distillation. It further introduces a modality-aware feature distillation that dynamically adjusts the distillation weights between question tokens and visual tokens based on their relative importance during training. MoInCL [113] proposes a pseudo-target generation module to synthesize input-target pairs for various task types. These pseudo samples are then mixed with current task data for knowledge distillation, effectively mitigating catastrophic forgetting. QUAD [114] leverages only past task questions for regularization to avoid overfitting to the current task’s answer space. It further introduces attention consistency distillation to preserve both intra-modal and inter-modal attention consistency across tasks.
4.3.2 Parameter-Level Regularization
Feature-level distillation strategies preserve prior knowledge by indirectly constraining model updates through additional losses, while parameter-level regularization, as depicted in Figure 5 (b), achieves the same goal by directly regularizing the model parameters. For MLLMs, some methods posit that only a small subset of parameters plays a critical role during downstream task training, while the majority contribute minimally. Consequently, continual learning can be facilitated by selectively protecting these important parameters to mitigate forgetting. For example, Model Tailor [115] first defines two metrics, a sensitivity-based score and a salience-based score, to capture a minimal set of parameters (model patch) that are important when fine-tuning the current downstream task. To enhance target task learning, it further applies a Hessian-based weight compensation method, termed the patch decorator. SPIDER [116] assesses the importance of generalizations in the original weights using weight magnitude analysis and evaluates the importance of specialized knowledge in the fine-tuned model through gradient changes during training. By comparing these two importance distributions, it adaptively fuses the original and fine-tuned weights, effectively preserving the model’s original generic capabilities while incorporating new task knowledge. SEFE [117] categorizes continual learning forgetting in MLLM into superficial forgetting and essential forgetting. To address superficial forgetting, it proposes the answer style diversification paradigm that standardizes response formats across tasks through style transformation. For essential forgetting, it introduces RegLoRA which aggregates positions marked by prior task regularization masks and optimizes them to approach zero, preserving parameters encoding critical prior knowledge.
4.4 Replay-based Approach
As illustrated in Figure 5 (c), replay techniques in MLLMs generally necessitate the storage of paired image inputs and corresponding instructional data or tokens. For example, VQACL [118] proposes a rehearsal-based method that disentangles sample-specific and sample-invariant features to separate reasoning skills and visual concepts. Sample-specific features are captured via self-attention, while sample-invariant features are learned through dynamically updated prototypes to balance plasticity and stability. SGP [119] leverages scene graphs as concise visual representations to replay pseudo scene graph question answer triplets, where a Symbolic Replay Model based on a language model captures symbolic reasoning mechanisms and learns task-specific mappings between scene graphs and QA pairs, while a Unified VQA Model is trained on both current and replayed samples to adapt to new tasks while retaining prior knowledge, addressing the challenges of multi-modal reasoning and privacy concerns in continual VQA. ProtoGroup [120] proposes a multi-prototype grouping framework, which clusters related prototypes into groups to extract more stable and accurate sample-invariant representations, and employs a memory sample selection method based on prototype group correlations to choose representative and memorable samples for memory updates. Lin et al. [121] leverage memory replay with TF-IDF and K-means clustering to store representative samples, while knowledge distillation guides the model through teacher-student learning to preserve prior knowledge. Additionally, task-specific projection layers are introduced to regularize feature embeddings, minimizing drift across tasks and ensuring continuity in learning, thereby addressing catastrophic forgetting in autonomous driving VQA. Adapt-$\infty$ [122] introduces a dynamic data selection framework that clusters samples into pseudo-skill groups via gradient-based representations, and adaptively selects high-impact samples using a multi-way scoring mechanism. To maintain efficiency, it permanently prunes semantically redundant data while preserving task balance. OASIS [123] introduces an adaptive sample selection framework for continual visual instruction tuning. It selects informative samples via forward-pass Fisher Information and reduces redundancy through gradient-based similarity. To ensure efficiency, it dynamically adjusts selection using exponential moving averages without backward computation.
4.5 Discussions
For continual learning tasks in Multimodal Large Language Models, approaches leveraging architectural extensions, particularly mixture-of-experts models, continue to dominate. However, although existing multimodal continual learning methods utilize multimodal datasets and architectures, they are primarily designed for general-purpose tasks. In other words, these methods do not differ substantially from continual learning strategies applied to language-only LLMs. Consequently, challenges related to modality-specific forgetting and the influence of inter-modal interactions during continual learning remain largely unexplored and warrant further investigation.
5 Continual Learning in Vision-Language Action Models
Vision-Language Action (VLA) [218, 219, 20, 220] models refer to exploiting large models to predict action and control behaviors of robots [221, 222] with the aid of a large vision language model. It is of great significance to replace the manipulation of humans in certain risky situations and therefore improve the level of automation. However, robots are inherently adapting to a particular field and inevitably encounter the problem of poor generalization [223] when dealing with the complex real world. Consequently, continual learning in vision language action [124] aims to equip robots with the ability to adapt and acquire new skills in dynamic and ever-changing real-world scenarios like humans do. It is a specific aspect of lifelong robotic learning [224, 225, 226, 227, 228] that employs large multimodal models for robot control and is crucial for robots to accumulate knowledge in novel environments while retaining the ability for known skills to mitigate catastrophic forgetting. Figure 6 illustrates the task formulation of visual-language action models within a continual learning framework.
Due to its wide applicability and potential, many articles have investigated the problem from many aspects based on the general strategy. In the following section, we first introduce the problem setup (Section 5.1) with common training objectives and benchmarks, then we will categorize vision language action methods into three detailed types: replay-based (Section 5.4), architecture-based (Section 5.2) and regularization-based (Section 5.3) approaches, which will be displayed in separate paragraphs.
TABLE III: Summary of continual learning methods for Vision-Language Action Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Tag: I = Image, T = Text, S = Sensor.
Methods Taxonomy Benchmark / Dataset Backbone Modality Environmental Setting Code
Arch Reg Rep
LIBERO [124] LIBERO ViT-T [210], ResNet-T [229], BERT I T Simulated Link
LOTUS [125] LIBERO DINOv2 [230], BERT I T Realistic & Simulated Link
QueST [126] LIBERO, MetaWorld [231] ResNet18, CLIP I T Simulated Link
LEGION [127] LEGION RoBERTa I T Realistic & Simulated Link
CAMA [128] ALFRED [232] ResNet, Bi-LSTM I T Simulated Link
M2Distill [129] LIBERO ResNet, BERT I T S Simulated -
iManip [130] RLBench [233] CLIP, U-Net I T Realistic & Simulated -
RWLA [131] LIBERO R3M [234], SS [235] I T Simulated -
5.1 Problem Setup
Lifelong learning [30, 236] in vision language action extends the concept of the traditional robot action control framework by requiring one robot to continually acquire and adapt across a sequence of tasks $\{T^{1},\dots,T^{K}\}$ over its lifespan while retaining knowledge of previous tasks. This continual learning problem for robots can be formulated as a finite-horizon Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},H,\mu_{0},R)$, where $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function, $H$ is the maximum horizon for each episode of a task, $\mu_{0}$ is the initial state distribution, and $R:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function. During the whole training procedure, each task $T^{k}$ trains the policy $\pi$ from the last task characterized by its own initial state distribution $\mu_{0}^{k}$, while $\mathcal{S},\mathcal{A},\mathcal{T}$ and $H$ remain unchanged across all tasks.
Training objectives. Continual learning in the vision-language action model sequentially trains one single policy $\pi$ through multiple tasks. For each task $T^{k},k\in\{1,\cdots,K\}$, the robot is trained on a dataset of $N$ demonstrations $D^{k}=\{\tau^{k}_{i}\}_{i=1}^{N}$ with a corresponding language description $l^{k}$ guiding the goal of the robot. Each trajectory $\tau^{k}_{i}$ in expert demonstrations with length $L^{k}<H$ consists of state-action pairs $\{(s_{t}^{k},a_{t}^{k})\}_{t=1}^{L^{k}}$ sampled from policy $\pi$ conditioned with task $T^{k}$. The policy is trained with a behavioral cloning loss [237] that aims to mimic the action demonstrations in the dataset:
$\min_{\pi}\mathcal{L}(\pi)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{s_{t}^{k},a_{t}^{k}\sim D^{k}}\left[\sum_{t=1}^{L^{k}}-\log\pi(a_{t}^{k}|s_{\leq t}^{k};T^{k})\right].$ (16)
In the scenario of continual learning, the single policy $\pi$ has to continually adapt to specific requirements of sequential tasks $T^{k}$, while retaining knowledge of gained skills to mitigate catastrophic forgetting.
Benchmarks. To evaluate the performance of robots in continual learning, many benchmarks have been constructed in continual learning scenarios from reinforcement learning to vision language action [238, 239, 124]. For example, Continual World [238] builds a continual reinforcement learning benchmark on robotic manipulation tasks. It provides a realistic, computationally accessible testbed that highlights the limitations of existing continual learning methods in reinforcement learning (RL) settings, which often struggle with forward transfer despite mitigating forgetting. LIBERO [124] builds an imitation learning environment for robots to analyze the continual learning problem of robots from many factors, such as algorithms, architectures, task sequences, variations, and so on, which provides a comprehensive environment for subsequent work.
Figure 6: Illustration of a continual learning framework applied to Vision-Language Action Models.
5.2 Architecture-based Approach
Architecture-based approaches investigate continual learning through structural design, with a focus on enhancing skill representation and reuse. One line of work explores skill-aware architectures that enable models to distinguish, merge, and update skills [240, 126, 125, 127], or construct extensible modular designs for skill composition [241, 242]. LOTUS [125] proposes a hierarchical imitation learning framework that incrementally builds a skill library via unsupervised segmentation of human demonstrations using open-vocabulary vision models. It updates or adds new skills through clustering and composes them via a meta-controller, achieving superior performance in lifelong manipulation through continual discovery and experience replay. IsCiL [243] introduces a prototype-based memory system that incrementally learns shareable skills for adaptation in dynamic environments. By mapping demonstrations to state embeddings and retrieving relevant skill prototypes, it employs adapter modules to preserve prior knowledge. LEGION [127] presents a lifelong reinforcement learning framework that combines Bayesian non-parametric clustering with language-conditioned embeddings. Using a Dirichlet Process Mixture Model (DPMM), it dynamically organizes skills and semantically aligns task understanding, enabling scalable skill accumulation without performance degradation. RLCM [244] formulates a modular lifelong RL framework that composes policies from learned neural modules. It adopts a two-phase learning process—online module selection and adaptation for new tasks, and offline consolidation via batch RL and experience replay—to retain and generalize skills over time.
Another line of research focuses on enhancing robotic action through architectural designs that leverage Large Language Models (LLMs) for high-level guidance in tasks such as planning [245] and skill demonstration [246, 247]. These approaches utilize pre-trained LLMs as knowledge bases, retrieving relevant information via textual prompts to facilitate skill acquisition and enable the transfer of prior knowledge to new tasks. Owing to their data efficiency and rapid inference capabilities, LLMs demonstrate strong potential across diverse robotic scenarios.
5.3 Regularization-based Approach
Regularization methods for continual learning usually try to keep important parameters or features unchanged and impose feature-level or parameter-level regularization during training, therefore maintaining the ability for past skills [248, 128, 129, 243]. CAMA [128] introduces Confidence-Aware Moving Average, a method that dynamically updates stored knowledge (logits) using confidence scores to balance past and current information without requiring task boundaries. PolyTask [248] proposes a method combining task-specific policy training with behavior distillation, which consolidates individual policies into a single model using offline data. With demonstration-guided reinforcement learning and distilling policy outputs, it handles continual actions and lifelong learning by avoiding interference between tasks. M2Distill [129] proposes a multi-modal distillation framework that preserves consistency across vision, language, and action modalities between incremental learning steps. It exploits: (1) multi-modal feature distillation to constrain latent space drift; and (2) policy distillation to maintain alignment of Gaussian mixture model (GMM) action distributions. The approach ensures skill retention without storing prior data, advancing lifelong learning for real-world robotic applications. ELIRL [249] introduces Lifelong Inverse Reinforcement Learning (Lifelong IRL), addressing the challenge of learning multiple sequential tasks and proposing Efficient Lifelong IRL to enable knowledge transfer across tasks, improving performance while minimizing user-provided data.
5.4 Replay-based Approach
Due to the spatial and temporal complexity of robotic manipulation in various and dynamic scenarios, it has become common practice to store part of previous samples and replay them [250, 38] during training procedure to retain the knowledge of previous tasks [130, 127, 241, 131]. iManip [130] addresses the challenge of catastrophic forgetting in robotic manipulation. Given that traditional methods fail to retain previously learned skills, ignoring temporal complexity and action complexity, the introduced framework integrates a temporal replay strategy to preserve task continuity and an extendable PerceiverIO [251] architecture with adaptable action prompts. Experimental results highlight the efficiency, scalability, and robustness across complex, long-horizon tasks. Decision-RWKV [241] exploits a novel sequence modeling framework that integrates the efficiency of the Receptance Weighted Key Value (RWKV) with experience replay mechanisms. It leverages the linear complexity and memory retention capabilities of RWKV to enable efficient lifelong learning across diverse manipulation tasks, and the experience replay combines a replay buffer for historical experiences reuse while acquiring new skills. RWLA [131] combines retrieval-based local adaptation with a selective weighting mechanism to restore forgotten skills. The method retrieves relevant past demonstrations from a memory buffer, identifies critical failure-prone segments, and prioritizes these segments during local fine-tuning, which allows robots to adapt continually while maintaining performance across diverse manipulation tasks.
5.5 Discussions
Continual learning in Vision-Language Action models increasingly adopts hybrid strategies that integrate architectural expansion, replay mechanisms, and regularization constraints. While early methods often apply these components independently, recent advances highlight the importance of balancing them to achieve effective continual learning. Specifically, architectural modifications enable task-specific adaptability, replay mechanisms help preserve cross-task consistency, and regularization mitigates catastrophic interference across modalities. However, key challenges remain, particularly in dynamically coordinating these strategies over time and addressing the tendency of action policy components to forget more rapidly than vision-language alignment modules.
6 Continual Learning in Diffusion Model
Latent Diffusion Models (LDMs) have emerged as a powerful tool in artificial intelligence, capable of generating photorealistic and semantically consistent images from user-defined textual prompts. Building on this capability, various Personalized Diffusion Models (PDMs) have been developed to generate images tailored to specific user input data (also known as concepts) and textual prompts, significantly enhancing interactivity. However, as shown in Figure 7, when user inputs and customization requirements arrive sequentially, these methods tend to forget previously learned concepts. Thus, overcoming forgetting in diffusion models under continual customization scenarios remains a key challenge.
In the following section, we first introduce the problem setup (Section 6.1), including training objectives and commonly used benchmarks. Then, we summarize existing research by categorizing it into architecture-based (Section 6.2) and regularization-based (Section 6.3) approaches.
6.1 Problem Setup
Training objectives. Diffusion model is a probabilistic generative model that learns the data distribution by gradually denoising samples drawn from a Gaussian distribution. Specifically, given an initial noise map $\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, a timestep $t\sim\text{Uniform}({1,\dots,T})$, and a user text prompt $p$, the generated image $x$ is obtained as $x=\mathcal{D}(\epsilon_{\theta}(c,t))$, where $c$ is the text embedding extracted from the prompt by a text encoder $\Gamma$ (i.e., $c=\Gamma(p)$). Formally, the diffusion model $\epsilon_{\theta}$ is trained to denoise the noisy image and predict the target as follows:
$\mathcal{L}_{\text{DM}}(\theta)=\mathbb{E}_{z,\varepsilon,c,t}[||\varepsilon-\epsilon_{\theta}(z_{t}|c,t)||^{2}],$ (17)
where $z_{t}$ is the noisy image at timestep $t$, and $\theta$ denotes the parameters of denoising model.
Figure 7: Illustration of a continual learning framework applied to Diffusion Models.
The above process outlines how to generate an image from random noise based on user prompts. To more faithfully capture the user’s intent, PDMs enhance generation by incorporating a small set of user-provided images alongside textual prompts, aiming to bind novel visual concepts (e.g., a specific person or object) to a unique identifier while preserving the model’s generalization ability. To achieve this, DreamBooth [252] introduces a class-specific prior preservation loss, which fine-tunes the model to generate the user-specified subject in various contexts while preventing overfitting and maintaining prior knowledge about the general class. The loss is formulated as:
$\mathcal{L}_{\text{PDM}}(\theta)$ $=\mathbb{E}_{z,\varepsilon,c,t}\left[\left\|\varepsilon-\epsilon_{\theta}(z_{t}|c,t)\right\|^{2}\right]$ (18)
$+\lambda\,\mathbb{E}_{z^{p},\varepsilon,c^{p},t}\left[\left\|\varepsilon-\epsilon_{\theta}(z_{t}^{p}|c^{p},t)\right\|^{2}\right],$
where $z_{t}$ denotes the latent code of an image with the new concept, and $z_{t}^{p}$ corresponds to a class-specific image used for prior preservation. $c$ and $c^{p}$ are the respective conditioning vectors; for instance, $c$ may be “a photo of [$\mathcal{V}^{*}$] dog”, while $c^{p}$ is “a photo of dog”. The special token [$\mathcal{V}^{*}$] is a unique identifier representing the user-specific concept learned from the provided images. $\lambda$ controls the regularization strength.
Continual learning in diffusion models aims to go beyond single-concept personalization by enabling the incremental acquisition of new concepts while avoiding catastrophic forgetting and eliminating the need to retrain on previous data. Specifically, as shown in Figure 7, assume a user intends to continuously generate content based on newly introduced specific concepts. The generation tasks are defined as $\mathcal{T}={\{\mathcal{T}^{k}}\}_{k=1}^{K}$ with corresponding datasets ${\{\mathcal{D}^{k}}\}_{k=1}^{K}$. Each $k$-th generation task consists of $n_{k}$ pairs of concept images $x_{i}^{k}$ and prompts $c_{i}^{k}$, where $K$ denotes the total number of tasks, and $n_{k}$ is typically in the range of 3 to 5 in accordance with existing studies. Accordingly, to learn the $k$-th personalized concept, the model is optimized using the following training loss:
$\mathcal{L}_{\text{CDM}}(\theta)$ $=\mathbb{E}_{z^{k},\varepsilon,c^{k},t}\left[\left\|\varepsilon-\epsilon_{\theta}^{k}(z_{t}^{k}|c^{k},t)\right\|^{2}\right]$ (19)
$+\lambda\,\mathbb{E}_{z^{k,p},\varepsilon,c^{k,p},t}\left[\left\|\varepsilon-\epsilon_{\theta}^{k}(z_{t}^{k,p}|c^{k,p},t)\right\|^{2}\right],$
where $(z_{t}^{k},c^{k})$ denote the image latent code at timestep $t$ and text embeddings corresponding to the $k$-th generation task, and $(z_{t}^{k,p},c^{k,p})$ represent the prior pairs generated by the pre-trained model for the same task.
Backbone. Stable Diffusion (SD-1.5) [22] is a high-resolution image synthesis model based on LDMs, offering a powerful and efficient generative framework. It operates by encoding images into a lower-dimensional latent space using a pre-trained autoencoder, significantly reducing computational costs during training and inference. The model then performs denoising in this latent space with a U-Net backbone guided by a text encoder, such as CLIP, enabling high-fidelity image generation conditioned on text prompts. In continual personalized diffusion settings, LoRA-based fine-tuning is typically applied to the U-Net backbone to efficiently adapt to limited training data.
Benchmark. CLoG [253] proposes two benchmarks: Label-conditioned and Concept-conditional, to evaluate the effectiveness of continual learning approaches in the image generation field. The label-conditioned setting requires the model to learn a sequence of generative tasks conditioned on label indices, while the concept-conditional setting focuses on synthesizing a series of personalized concepts, denoted as $\mathcal{V}^{*}_{k}$, where $k$ indicates the $k$-th concept.
6.2 Architecture-based Approach
C-LoRA [132] introduces a continual concept learning setting within the LDM framework, where each user-defined concept is assigned a dedicated LoRA module injected into the self-attention keys and values. During training, each concept is represented by a unique placeholder token (e.g., [$\mathcal{V}^{*}$]) inserted into the prompt to associate the visual concept with a textual identifier. To mitigate catastrophic forgetting, C-LoRA freezes the LoRA parameters associated with previously learned concepts during new concept learning. Furthermore, it incorporates a self-regularization loss to penalize updates to weight regions that were modified by earlier concepts, thus preserving prior knowledge.
6.3 Regularization-based Approach
MuseumMaker [133] aims to retain the knowledge of past learned styles while incorporating new customized styles. To mitigate catastrophic overfitting and maintain the integrity of original concepts in diffusion models, it introduces a style distillation loss that captures disentangled style representations. Furthermore, it proposes a dual regularization scheme for LoRA training, which constrains model optimization from both the weight and feature perspectives to achieve balanced knowledge integration. LFS-Diffusion [134] proposes a data-free knowledge distillation strategy applied across diffusion time steps during training, aiming to constrain the model from overfitting to current personalized concepts and thereby preserve the original generative capabilities. During inference, an in-context generation mechanism is introduced to mitigate forgetting of previously learned knowledge and ensure better retention of past concepts. CIDM [135] introduces a concept consolidation loss that employs learnable layer-wise concept tokens and an orthogonal subspace regularizer to disentangle task-specific and task-shared knowledge. During inference, an elastic weight aggregation module merges low-rank weights of previously learned concepts based on their importance. To alleviate concept neglect in multi-concept composition, CIDM further proposes a context-controllable synthesis strategy that enhances regional feature expressiveness via layer-wise textual embeddings and guides generation through region-aware noise estimation. ConceptGuard [136] introduces a set of lightweight yet effective strategies to mitigate forgetting in continual concept learning for diffusion models. It employs shift embeddings for dynamic concept updates, concept-binding prompts to disentangle and compose concepts over time, and memory preservation regularization to slow down harmful parameter drift. Additionally, a priority queue adaptively manages concept replay based on their importance and temporal order. Jha et al. [137] utilize diffusion classifier scores as implicit class-conditional density estimators. This method simultaneously regularizes both parameter-space and function-space consolidation, enabling effective adaptation to evolving user concepts.
6.4 Replay-based Approach
Since each stage in the continual personalized diffusion setting typically involves only a limited number of training samples (e.g., 3-5 images), a feature replay strategy is commonly adopted. Specifically, $\text{L}^{2}\text{DM}$ [138] addresses the continual concept learning challenge by introducing the Rainbow-Memory Bank Strategy, which consists of a short-term memory bank for encoding current task features, a long-term memory bank for preserving historical knowledge, and a scoring function to balance knowledge retention and transfer. To further mitigate the forgetting of personalized details, $\text{L}^{2}\text{DM}$ proposes Elastic Concept Distillation, which enhances model performance by distilling knowledge between the current model and the previous task model, using features retrieved from the short-term memory bank as input to both.
6.5 Discussions
In this section, we discuss continual learning in diffusion models, with a particular focus on continual personalized generation. Due to the limited training data per task and the relatively low variance across tasks in this setting, regularization-based approaches remain predominant. Nevertheless, we argue that a major limitation in this area lies in the absence of standardized evaluation benchmarks and robust quantitative metrics, which are essential for fairly assessing different methods. Furthermore, exploring broader continual learning scenarios within diffusion models represents a promising and underexplored research direction.
7 Future Directions
Despite significant advances in continual learning for generative models, critical challenges persist across multiple dimensions. In this section, we further identify promising research directions at the intersection of continual learning and generative modeling, aiming to stimulate further progress in the field.
Efficient Continual Learning Mechanisms. Current mainstream continual learning approaches often rely on model expansion to retain task-specific knowledge, which leads to substantial storage and computational overhead during deployment. These challenges are particularly pronounced for large-scale generative models, where the large number of parameters causes latency and resource limitations. Consequently, there is increasing interest in enhancing efficiency through techniques such as model compression [254, 255, 256], pruning [257, 258, 259, 260, 261, 262], and quantization [263, 264, 265]. In addition to structural optimization, recent progress in dataset compression and distillation [266, 267, 268, 269] highlights the potential of compressing replay data as a promising direction. A key idea is to encode as much task-relevant information as possible from previous data or features into a small number of representative tokens. This strategy not only helps mitigate catastrophic forgetting but also reduces the risk of data privacy leakage by avoiding the storage of raw data. Overall, integrating continual learning with efficient model and data compression techniques provides a compelling path toward improving the practicality and scalability of generative models in real-world applications.
Learning Paradigm for Continual Learning. While conventional continual learning methods predominantly rely on supervised fine-tuning (SFT) for incremental knowledge acquisition, emerging reinforcement learning (RL)-based training paradigms, particularly exemplified by GRPO [270], have demonstrated stronger generalization capabilities across a variety of model architectures. This performance advantage has been consistently observed not only in large language models [271, 272, 273] but also in multimodal large language models [274, 275, 276, 277] and vision-language action models [278, 279], suggesting that RL-based methods may serve as a more robust alternative to traditional SFT strategies in continual learning settings. Recent studies [280] have revealed a fundamental difference between the two paradigms: SFT tends to overfit specific training distributions through pattern memorization, whereas RL emphasizes structural comprehension of task objectives, enabling better generalization to novel scenarios. This distinction provides strong motivation for exploring continual learning within an RL framework, which may naturally support long-term adaptation, task abstraction, and reinforcement-driven retention strategies.
Self-generation for Continual Learning. A core challenge in continual learning lies in the inaccessibility of previous task data. In discriminative models, a widely adopted solution involves training auxiliary generative models (e.g.,, GANs [281] or diffusion models) to synthesize pseudo-samples of prior tasks, thereby mitigating catastrophic forgetting [282, 283, 284, 285]. While effective, this approach introduces significant computational overhead and may suffer from a distributional mismatch between generated and original data. In contrast, generative models offer a more integrated solution by leveraging their inherent capacity to reconstruct and regenerate data. These models can autonomously replay past task distributions during the learning of new tasks, without the need for separate generative modules. This capability offers two major benefits: reducing additional architectural complexity and achieving better alignment with the original data distribution. As a result, self-generation in generative models presents a promising direction for building more efficient, scalable, and distributionally consistent continual learning systems.
Continual Learning in Large Scale Models. Most existing studies on continual learning for generative models have focused on small- to medium-scale architectures (e.g., models with up to 7B and 13B parameters). However, emerging evidence [286, 287] suggests that continual learning behaviors can differ substantially as the model scale increases. This observation highlights two critical research directions. First, there is a pressing need to establish specialized evaluation benchmarks that mitigate information leakage risks [288], particularly given the large and often proprietary datasets used to train large-scale models. Second, it is essential to explore parameter-efficient knowledge integration strategies—such as Chain-of-Thought reasoning (CoT)[166, 289], In-Context Learning (ICL)[290, 291, 292], Agents [293, 294, 295], and Retrieval-Augmented Generation (RAG) [296, 297]—as conventional fine-tuning becomes increasingly infeasible for models with extensive parameter counts.
Continual Learning for Multimodal Generation Beyond Vision and Language. While most existing research in continual generative learning focuses on vision-language tasks, real-world multimodal systems are increasingly incorporating richer modalities such as audio [298], 3D scenes [299], and motion data [300]. In such settings, modality-continual learning becomes a critical yet underexplored challenge. For example, a pre-trained MLLM may be required to incorporate speech input or generate spoken responses. Without appropriate continual learning mechanisms, learning the new audio modality could result in severe forgetting of previously acquired vision-language capabilities. Similarly, in VLA models deployed in robotics or embodied agents, the integration of new sensor modalities (e.g., depth and haptic signals) necessitates stable expansion of generative capacity without overwriting existing knowledge. These scenarios highlight the need for continual learning methods capable of integrating new modalities while preserving alignment and generative quality across previously learned ones.
Toward Unified Optimization Frameworks for Continual Learning in Generative Models. Although generative models vary widely in data modalities and task objectives, increasing architectural convergence is fostering the development of unified continual learning frameworks. For example, both multimodal large language models and vision-language action models frequently rely on large language models to enable cross-modal reasoning. Furthermore, recent attempts to integrate diffusion models with MLLMs [301, 302, 303] underscore a broader trend toward unifying generation and understanding within generative AI. These shared characteristics suggest a promising direction: designing continual learning strategies that generalize across generative model families. Unifying optimization across modalities and architectures may mark a critical step toward deploying generative AI in complex, dynamic real-world scenarios.
8 Conclusion
In this survey, we have reviewed recent advances in continual learning for generative models, providing a systematic overview across four major categories: Large Language Models, Multimodal Large Language Models, Vision-Language Action Models, and Diffusion Models. We further unify these models by identifying shared methodological dimensions that underpin their continual learning paradigms. Specifically, we formalize the continual learning problem in each domain, introduce relevant datasets and model architectures, and categorize representative approaches within each subfield. We hope that this comprehensive and in-depth survey, along with the accompanying real-time updated GitHub repository, will not only consolidate the current understanding of continual learning in generative modeling but also serve as a valuable reference and catalyst for future research.